# Urban4D: Semantic-Guided 4D Gaussian Splatting for Urban Scene Reconstruction

### v1

@@ 3.5 Optimization Strategy
$\mathcal{L}_{\text{total}}=$ $\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{sem}}$
$+\lambda_{4}\mathcal{L}_{\text{ground}}+\lambda_{5}\mathcal{L}_{\text{depth}}+\lambda_{6}\mathcal{L}_{\text{sky}}$ (9)
where $\{\lambda_{i}\}_{i=1}^{6}$ are weighting coefficients balancing different loss terms. Each loss term serves a specific purpose in our optimization.
@@
Ground Consistency. As detailed in Sec. 3.4, the ground consistency loss enforces local geometric similarity as defined in Eq. 8.
@@
Figure 5: comparison of reconstruction quality across consecutive frames. DeformGS [40] struggles significantly with reconstructing dynamic objects, resulting in severe artifacts and a failure to accurately represent motion. PVG [7] captures dynamic vehicles to some extent but suffers from noticeable blurring, particularly in the lower parts of the objects. In contrast, Urban4D delivers superior reconstruction quality, maintaining high fidelity and preserving clear details throughout the dynamic objects.
Table 1: Quantitative comparison with existing methods on the nuScenes [4] dataset. The best results in each group are highlighted in bold. Our method outperforms previous approaches across all evaluation metrics.
Method Full Image Non-Sky Human Vehicle Dynamic
@@
EmerNeRF [38] 26.12 0.830 26.50 0.840 23.45 0.733 24.69 0.808 24.12 0.767
PVG [7] 26.23 0.834 26.72 0.841 23.98 0.743 24.73 0.815 24.33 0.774
Urban4D (Ours) 26.47 0.839 28.93 0.878 24.88 0.783 26.05 0.855 25.33 0.810
@@ 4.1 Implementation Details
We initialize Gaussian points from both LiDAR points (with projected RGB and semantic values) and 200K random points sampled within a sphere. Following previous 3DGS-based methods to predict reliable semantic Gaussians, we use SegFormer [35] as our pre-trained segmentation model. The learning rate of MLP starts from $1.6\times 10^{-4}$ and decreases to $1.6\times 10^{-6}$. For each loss term, the weighting coefficients are empirically set to $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, $\lambda_{3}=0.01$, $\lambda_{4}=0.0001$, $\lambda_{5}=0.1$ and $\lambda_{6}=0.01$. All experiments are conducted on a single NVIDIA V100s with 32GB memory.
@@ 4.2 Datasets
Our experiments are conducted on two popular autonomous driving datasets: PandaSet [7] and nuScenes [4]. We primarily evaluate our method on nuScenes [4], which consists of 1000 scenes captured in Boston and Singapore under diverse urban scenarios with multimodal sensor data, featuring challenging weather conditions and complex urban scenes. We perform ablation studies on PandaSet [7], which is a comprehensive autonomous driving dataset captured in San Francisco, containing 103 sequences of synchronized LiDAR-camera data and images. For both datasets, we use the officially provided camera poses.
@@ 4.3 Results and Comparisons
Results on nuScenes [4]. We evaluate our method against previous approaches on the nuScenes [4] dataset, with a comprehensive comparison across different scene components and object categories. As shown in Table 1, our method demonstrates consistent improvements over NeRF-based methods [30, 38], traditional 3DGS [13], PVG [7] and DeformGS [40] across all evaluation metrics, with significant improvements in dynamic object reconstruction.
For full image reconstruction, our method achieves 26.47 PSNR and 0.839 SSIM, surpassing PVG [7] by 0.24 PSNR. The improvement is more pronounced in non-sky regions, where our approach shows a substantial gain of 2.21 PSNR (28.93 vs 26.72) and 0.037 SSIM (0.878 vs 0.841) compared to PVG [7], indicating better handling of complex geometric structures in urban environments.
The advantages of our method are most evident in dynamic object reconstruction. For human instances, we achieve 24.88 PSNR and 0.783 SSIM, outperforming PVG [7] by 0.90 PSNR and 0.040 SSIM. In vehicle reconstruction, our method demonstrates even larger improvements, reaching 26.05 PSNR and 0.855 SSIM, surpassing PVG [7] by 1.32 PSNR and 0.040 SSIM. These significant gains can be attributed to our dynamic-focused semantic decomposition strategy, which enables specialized treatment of moving objects.
We present qualitative comparisons in Fig. 5. For dynamic objects like vehicles and pedestrians, our method shows significant improvements over PVG [7] and DeformGS [40], with notably reduced motion blur and sharper object boundaries. This is particularly evident in regions with fast-moving objects, where previous methods tend to produce blurry reconstructions. Meanwhile, our approach also demonstrates better reconstruction quality for static scene elements such as roads, preserving more detailed textures and geometric structures. These visual improvements align well with our quantitative results, where we achieve consistently higher scores across both dynamic and static regions. The qualitative results in Fig. 5 further support these findings. In consecutive frame reconstruction, DeformGS [40] fails to handle dynamic objects, producing severe artifacts. While PVG [7] captures the overall shape of moving vehicles, it suffers from noticeable blurring artifacts, particularly in the lower parts of the vehicles. In contrast, our method achieves clearer and more consistent reconstruction across all dynamic objects.
Table 2: Quantitative comparison with state-of-the-art methods on the PandaSet [34] dataset. We report image reconstruction and novel view synthesis metrics. The best results are shown in bold.
@@
Urban4D (Ours) 27.74 0.850 26.56 0.814
Results on PandaSet [34]. To further validate the generalization capability of our method, we conduct additional experiments on the PandaSet [34] dataset. As shown in Table 2, our approach demonstrates consistently superior performance across both image reconstruction and novel view synthesis tasks. For image reconstruction, our method achieves 27.74 PSNR and 0.850 SSIM, outperforming PVG [7] by 0.59 PSNR and 0.014 SSIM. The performance advantage is even more pronounced when compared to vanilla 3D Gaussian Splatting (3DGS [13]), showing improvements of 4.07 PSNR and 0.107 SSIM. This significant gap highlights the effectiveness of our dynamic-focused scene decomposition strategy.
In the more challenging novel view synthesis task, our method maintains its lead with 26.56 PSNR and 0.814 SSIM, surpassing PVG [7] by 0.64 PSNR and 0.016 SSIM. Notably, while all methods experience some performance degradation in novel view synthesis compared to image reconstruction, our method shows the smallest drop (1.18 PSNR vs 1.23 PSNR for PVG [7]), indicating better scene understanding and geometry modeling. The consistent performance improvements across different datasets and evaluation metrics demonstrate the robustness and generalizability of our approach across varied urban scenes and capture conditions.
@@ 4.4 Ablation Study
Effectiveness of consistency regularization. To validate the effectiveness of our ground surface consistency module, we conduct extensive experiments on the PandaSet [34] dataset. Table 3 shows the quantitative results of different variants of our method. The baseline model without road consistency achieves 26.31 PSNR and 0.810 SSIM. After incorporating our consistency constraint, the performance significantly improves by 0.25 dB in PSNR (26.56 vs. 26.31) and 0.004 in SSIM (0.814 vs. 0.810). This demonstrates that enforcing local geometric and appearance consistency on ground surfaces effectively enhances the reconstruction quality.
We further investigate the impact of different numbers of nearest neighbors in our module. As shown in Table 3, using 16 neighbors yields the best performance (26.56 PSNR), outperforming both smaller (k=8, 26.45 PSNR) and larger (k=32, 26.48 PSNR) neighbor sizes. We hypothesize that 8 neighbors are insufficient to capture the local surface structure, while 32 neighbors may introduce noise from points that are too far away. The optimal choice of 16 neighbors provides a good local structure preservation.
Table 3: Ablation study on road consistency module. We evaluate different design choices and parameter settings of our road consistency constraint on PandaSet [7].
@@
Baseline (w/o road consistency) 26.31 0.810
KNN-8 26.45 0.812
KNN-16 (Ours) 26.56 0.814
KNN-32 26.48 0.812
Table 4: Ablation study on time embedding. We report PSNR and SSIM on the PandaSet [34].
[…]
w/o Time Embeddings 26.40 0.810
w/ Time Embeddings 26.56 0.814
Effectiveness of Time Embeddings. We also validate the effectiveness of our temporal modeling strategy through ablation studies. As shown in Table 4, equipping each Gaussian with a learnable time embedding improves the rendering quality (PSNR: 26.40 → 26.56, SSIM: 0.810 → 0.814) on the PandaSet [34] dataset. This demonstrates that point-specific temporal context helps capture more accurate deformation patterns compared to directly mapping spatial positions to deformations.
@@ 5 Conclusions
In conclusion, Urban4D provides a novel semantic-guided decomposition and 4D representation strategy for reconstructing dynamic urban scenes. By leveraging 2D semantic information, our approach effectively separates static and potentially dynamic components. We introduced specialized processing methods for different elements: a 4D Gaussian Splatting representation for dynamic objects and KNN-based consistency regularization for the ground surface. Both quality and quantitative results demonstrated that Urban4D improves rendering quality across various scene components, particularly for dynamic objects and low-texture regions.

### v2

@@ 3.5 Optimization Strategy
$\mathcal{L}=$ $\lambda_{1}\mathcal{L}_{\text{L1}}+\lambda_{2}\mathcal{L}_{\text{SSIM}}+\lambda_{3}\mathcal{L}_{\text{sem}}$
$+\lambda_{4}\mathcal{L}_{\text{static}}+\lambda_{5}\mathcal{L}_{\text{ground}}+\lambda_{6}\mathcal{L}_{\text{depth}}+\lambda_{7}\mathcal{L}_{\text{sky}},$ (11)
where $\{\lambda_{i}\}_{i=1}^{7}$ are weighting coefficients balancing different loss terms. Each loss term serves a specific purpose in our optimization.
@@
Static invariance & Ground Consistency. As detailed in Sec. 3.3.
@@
Figure 4: Comparison of reconstruction quality across consecutive frames. DeformGS [46] struggles significantly with reconstructing dynamic objects, resulting in severe artifacts and a failure to accurately represent motion. PVG [7] captures dynamic vehicles to some extent but suffers from noticeable blurring, particularly in the lower parts of the objects. In contrast, UrbanGS delivers superior reconstruction quality, maintaining high fidelity and preserving clear details throughout the dynamic objects.
Table 1: Quantitative comparison with existing methods on the nuScenes [4] dataset. D: DINO features, F: Optical Flow, S: Semantic map, B: Bounding box. * indicates methods that use ground-truth (GT) bounding boxes.
Method Extra Inputs Full Image Non-Sky Human Vehicle Dynamic
@@
EmerNeRF [44] S 26.12 0.830 26.50 0.840 23.45 0.733 24.69 0.808 24.12 0.767
PVG [7] S 26.23 0.834 26.72 0.841 23.98 0.743 24.73 0.815 24.33 0.774
streetGS [43] S+B 25.46 0.831 25.50 0.820 22.54 0.659 25.87 0.837 24.17 0.771
4dgf [10] B 21.20 0.694 24.37 0.769 20.36 0.619 22.74 0.762 22.39 0.698
OmniRe [8] S+B 26.41 0.837 26.73 0.845 23.71 0.737 25.95 0.856 25.24 0.805
UrbanGS (Ours) S 26.92 0.848 27.61 0.861 24.92 0.767 26.01 0.838 25.43 0.818
streetGS [43] S+$\mathrm{B}^{*}$ 25.89 0.845 26.01 0.858 22.83 0.705 26.88 0.852 25.10 0.803
4dgf [10] $\mathrm{B}^{*}$ 27.48 0.852 27.81 0.865 25.16 0.789 27.14 0.858 26.22 0.843
OmniRe [8] S+$\mathrm{B}^{*}$ 29.15 0.873 29.54 0.886 26.10 0.835 27.23 0.861 26.68 0.856
@@ 4.1 Implementation Details
We initialize Gaussian points from both LiDAR points (with projected RGB and semantic values) and 200K random points sampled within a sphere. Following previous 3DGS-based methods to predict reliable semantic Gaussians, we use SegFormer [41] as our pre-trained segmentation model. Our approach builds upon PVG [7], with all parameters configured identically to its original implementation. The learning rate of MLP starts from $1.6\times 10^{-4}$ and decreases to $1.6\times 10^{-6}$. For each loss term, the weighting coefficients are empirically set to $\lambda_{1}=0.8$, $\lambda_{2}=0.2$, $\lambda_{3}=0.01$,$\lambda_{4}=0.01$, $\lambda_{5}=0.0001$, $\lambda_{6}=0.1$ and $\lambda_{7}=0.01$. All experiments are conducted on a single NVIDIA V100s.
@@ 4.2 Datasets
Our experiments are conducted on two widely-used autonomous driving datasets: nuScenes [4] and PandaSet [7]. The nuScenes dataset [4] comprises 1,000 scenes captured in Boston and Singapore under diverse urban scenarios. PandaSet [7] is a comprehensive dataset collected in San Francisco, containing 103 sequences with synchronized LiDAR and camera data. Both datasets provide ground-truth (GT) 3D bounding boxes. To ensure fair comparisons, we also utilize a pretrained state-of-the-art 3D multi-object tracking model, MCTrack [37], to generate predicted 3D bounding boxes for methods requiring such inputs. Additionally, we conducted experiments on the Waymo dataset [30]; detailed results and analyses are available in the supplementary material.
@@ 4.3 Results and Comparisons
Results on nuScenes [4]. We comprehensively evaluate our method against previous state-of-the-art approaches on the nuScenes [4] dataset, reporting quantitative results in Table 1. Our approach achieves superior performance among methods that do not rely on ground-truth (GT) bounding boxes, surpassing recent semantic-based methods (e.g., PVG [7], EmerNeRF [44]) as well as bounding-box-based approaches utilizing predicted boxes (e.g., streetGS [43], OmniRe [8], 4dgf [10]). Specifically, for full-image reconstruction, our method achieves 26.92 PSNR and 0.848 SSIM, significantly outperforming PVG [7] by 0.69 PSNR and 0.014 SSIM. In non-sky regions, our gains become even more pronounced, reaching 27.61 PSNR and 0.861 SSIM, clearly exceeding PVG by 0.89 PSNR and 0.020 SSIM. Such improvements highlight our method’s effectiveness in accurately reconstructing detailed urban structures.
Remarkably, our approach not only outperforms all methods relying on predicted bounding boxes, but also surpasses the streetGS [43] method even when it leverages GT bounding boxes. This result demonstrates the strong robustness and effectiveness of our method, as we achieve superior accuracy without explicit bounding-box supervision. Additionally, methods that rely heavily on bounding boxes experience significant performance degradation when switching from GT to predicted bounding boxes; notably, the 4dgf [10] method struggles significantly under predicted bounding-box inputs, failing to converge and exhibiting very low performance.
In dynamic object reconstruction, our method further demonstrates clear advantages. For human instances, we achieve 24.92 PSNR and 0.767 SSIM, substantially outperforming PVG by 0.94 PSNR and 0.024 SSIM. For vehicle instances, we achieve 26.01 PSNR and 0.838 SSIM, representing a significant gain of 1.28 PSNR and 0.023 SSIM compared to PVG. These improvements underscore our model’s strength in effectively reconstructing dynamic content in challenging urban environments, without relying on external bounding-box supervision.
We present qualitative comparisons in Fig. 1. For dynamic objects like vehicles and pedestrians, our method shows significant improvements over PVG [7] and DeformGS [46], with notably reduced motion blur. Meanwhile, our approach also demonstrates better reconstruction quality for static scene elements such as roads, preserving more detailed textures and geometric structures. These visual improvements align well with our quantitative results, where we achieve consistently higher scores across both dynamic and static regions. The qualitative results in Fig. 5 further support these findings. In consecutive frame reconstruction, DeformGS [46] fails to handle dynamic objects, producing severe artifacts. While PVG [7] captures the overall shape of moving vehicles, it suffers from noticeable blurring artifacts, particularly in the lower parts of the vehicles. In contrast, our method achieves clearer and more consistent reconstruction across all dynamic objects.
Table 2: Quantitative comparison with state-of-the-art methods on the PandaSet [40] dataset. We report image reconstruction and novel view synthesis metrics. * indicates using ground-truth data
@@
OmniRe [8] 26.94 0.840 25.75 0.804
UrbanGS (Ours) 28.03 0.858 26.76 0.821
$\mathrm{OmniRe}^{*}$ [8] 29.02 0.882 27.73 0.855
Results on PandaSet [40]. We further evaluate our method on the PandaSet dataset [40], comparing it against recent state-of-the-art methods in both image reconstruction and novel view synthesis tasks (Table 2). Our approach consistently outperforms previous methods that rely on predicted bounding boxes or do not use bounding boxes at all, including OmniRe [8], PVG [7], EmerNeRF [44], and 3DGS [17]. Although methods utilizing ground-truth bounding boxes achieve higher performance, our method demonstrates superior robustness and effectiveness under realistic conditions (i.e., without ground-truth bounding boxes). These results highlight the effectiveness of our proposed approach in accurately reconstructing urban scenes. For additional qualitative results, please refer to the Supplementary.
@@ 4.4 Ablation Study
We conduct an ablation study to evaluate the contribution of each module in our method on the PandaSet [7]. The results are summarized in Table 3.
Effect of Static Invariance. The absence of the static invariance module results in a notable decline in performance, with PSNR and SSIM dropping to 26.56 and 0.814, respectively. This underscores the importance of incorporating static priors to improve the reconstruction of static regions, thereby ensuring more robust and accurate results.
Effect of Road Consistency. Removing the road consistency module causes a performance drop. This demonstrates that enforcing road consistency is essential for preserving the geometric and textural integrity of roads, which significantly enhances the overall reconstruction quality.
Effect of 4D Representation. Excluding the 4D representation leads to the most significant performance degradation, with PSNR reducing to 26.43 and SSIM dropping to 0.810. This highlights the critical contribution of the 4D representation in modeling potentially dynamic objects and handling temporal variations, which are essential for reconstructing complex urban scenes.
Visualization of Static Regularization. Figure 5 presents a visual comparison of the effect of static regularization on Waymo [30]. When the Static Invariance and Road Consistency modules are removed, the reconstructed ground becomes significantly more blurred. In contrast, incorporating these modules results in much clearer and more visually pleasing reconstructions, demonstrating their effectiveness.
Table 3: Ablation study on each module. We evaluate the effect of each module on PandaSet [7].
@@
Baseline (w/o static invariance) 26.56 0.814
Baseline (w/o road consistency) 26.60 0.816
Baseline (w/o 4D Representation) 26.43 0.810
UrbanGS(Ours) 26.76 0.821
Figure 5: Ablation study on the effectiveness of static regularization. Results without static regularization (left) are blurry, while adding it (right) produces sharper details.
@@ 5 Conclusions
In conclusion, UrbanGS provides a novel semantic-guided decomposition strategy for reconstructing urban scenes. By leveraging 2D semantic information, our approach effectively separates static elements from potentially dynamic components. We introduced specialized processing methods for different elements: A static Invariance module to leverage static priors to enhance the reconstruction of stationary regions and KNN-based consistency regularization for the ground surface. And a 4D Gaussian Splatting representation for potentially dynamic objects. Both qualitative and quantitative results demonstrated that UrbanGS improves rendering quality across various scene components.